OpenStudy (anonymous):

The area under f(x) = 2x – 3 on the domain 3 ≤ x ≤ 7 is?

OpenStudy (anonymous):

Do not tell me the answer, I know the answer. I just want to know how to do the problem.

OpenStudy (anonymous):

integrate the function over the specified interval

OpenStudy (anonymous):

\[\int\limits_{3}^{7}(2x-3)\,dx\]

OpenStudy (anonymous):

i'm assuming you're in calculus. are you?

OpenStudy (anonymous):

AP calculus :| Something I very much regret.

OpenStudy (anonymous):

oh, come on now... you'll do great. do your best to keep a good attitude .

OpenStudy (anonymous):

Lol, I try. Passed the first semester with a B-, so I'll manage somehow. Anyways, is there some type of formula to go by?

OpenStudy (anonymous):

do you know how to integrate that?

OpenStudy (anonymous):

Not exactly

OpenStudy (anonymous):

okay, i'll help you with this and then i'll pass along some "cheat sheets" which will hopefully make your life a little better! ;) first off, we're gonna break this up because the function is a sum (or difference). \[\int\limits_{3}^{7}(2x-3)\,dx = \int\limits_{3}^{7}2x\,dx-\int\limits_{3}^{7}3\,dx = \left[\frac{2x^2}{2}-3x\right]_{3}^{7}\]

OpenStudy (anonymous):

Then what happens...? Lol. I'm sorry, I'm very slow when it comes to math so please be patient with me.

OpenStudy (anonymous):

evaluate at the points... \[(7^2-3(7))-(3^2-3(3))\]

OpenStudy (anonymous):

(49 - 21) - (9 - 9) = 28

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

check out this link. it's where i go the sheets from. http://tutorial.math.lamar.edu/Classes/CalcI/CalcI.aspx

OpenStudy (anonymous):

These sheets will come in handy. Appreciate your help. :)

OpenStudy (anonymous):

no worries!